OpenStudy (anonymous):

6^7/6^2 divide and simplify

OpenStudy (anonymous):

6^5 = 6*6*6*6*6 =?

OpenStudy (lgbasallote):

a^m/a^n = a^(m-n)

OpenStudy (anonymous):

so, here, it is 6^5=6*6*6*6*6=7776

OpenStudy (anonymous):

6^5 is the answer

Directrix (directrix):